[Brahms-l] Fw: SPY news : in the physics mode of running

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Sat Jan 03 2004 - 18:15:12 EST
FYI



====== Scheduling Physicist -YIP (ie. SPY) news channel ========

Hello,

(1) Status : 

We declared the beginning of physics mode running on Dec. 31, 2003
which was about a week earlier than previously scheduled.   Some time
after 14:30 that day, we had our first "physics" store with successful
rebucketing with 56x56 bunches.  Since then, though there have been
minor problems here and there, the machine has performed quite
well.  Though I shouldn't try to be too early to be too "gung ho", people
start to think that meeting our luminosity target is not too difficult ! :-)

I understand that STAR is eagerly using the collision data to evaluate 
whether their TPC can operate well in spite of the minor defect (a 
shorted circuit somewhere).   I was called this afternoon and told
that PHOBOS hasn't turned on their silicon detectors because they 
want to make sure that the re-trigger mechanism for handling the RHIC
abort kicker pre-fire (~ "misfire" to me) is working properly (otherwise, their 
silicon detectors could be damaged).   I did something to make the relevant
parties talk to each other.  I hope PHOBOS will be re-convinced soon 
(probably after the next ~2 successful beam dumps).


(2) The usual Time meeting will be held next Tuesday Jan. 6, 2004
at 13:30 in the Snyder Seminar Room (911A).  On Monday (Jan. 5), 
we will decide upon whether the access on Jan. 7, 2004 should be
materialized.

(3) Last Tuesday morning in my special & self-centered perspective ...

If you don't know, Tuesday morning, we were kind of a few steps away from 
opening a vacuum valve/section which might result in ~a week's downtime.
Because people suspected that the vacuum valve was closed and thus
blocked the beam from circulating into the yellow ring.  At the end, after
12 hours (?!), the smart machine specialist Greg found out that it was due 
to 4 magnets not having the correct currents.  Their statuses were mistakenly 
shown to be "ON" in a panel --- which was in turn due to a Programmable Logic 
Controller having a communication problem with some of its attached equipment
after the power dip on Sunday evening !

Greg went to look at a couple different pages which showed the magnet
currents read back ... and thus discovered the problem.   Believe it or not, 
in that exact morning, I actually looked at one of those panels myself 
but I didn't actually understand what those "zeroes" etc. meant.  I should
have asked my (n+1)th stupid question --- see ! again, there is really
NO stupid question.

I remember one evening a week or 2 ago, when I was in the Main Control 
Room, a similar situation occurred (though not in the same location),
ie., one beam going through but not the other.   After 30-45 minutes, people
actually realized that the current read back for a trimming magnet was much
lower than what it supposed to be.

Kin

====== ======= ====== ======= ====== ======= ====== =======

This SPY news channel solicits YOUR input of any sorts.
You'll be paid back with loads of gratefulness 

Let me know if you want to get off this private email mailing list.






_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l



smile_n.gif
Received on Sat Jan 3 18:07:22 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 03 2004 - 18:07:32 EST